OpenStudy (anonymous):

Find the standard equation for the ellipse using the characterstics given vertices: (0, PM7) foci (0, Sqrt PM 33)

OpenStudy (slaaibak):

Standard form of the ellipse: x^2/a^2 + y^2/b^2 = 1 and/or x^2/b^2 + y^2/a^2 = 1 since the focii and vertices are at the (0, k) we use: x^2/b + y^2/a = 1 c=foci = sqrt(33) a = 7 c^2 = a^2 - b^2 33 = 49 - b^2 b=4 So equation is: x^2/4^2 + y^2/7^2 = 1 x^2/16 + y^2/49 = 1

OpenStudy (slaaibak):

Third line in the message above should be: x^2/b^2 + y^2/a^2 = 1

OpenStudy (anonymous):

thanks so much again.